OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 12 and vertex at (11, -2). y = 2x2 - 44x + 120 y = 2x2 - 44x - 240 y = 2x2 - 44x + 240 y = -2x2 - 44x + 240 im thinking its either the last one or second the the last one.

OpenStudy (welshfella):

vertex form is a(x - h)^2 + k where (h,k) are the coordinates the the vertex

OpenStudy (welshfella):

and a is a number to be found

OpenStudy (anonymous):

how do i find it?

OpenStudy (welshfella):

so here h = 11 and k = -2

OpenStudy (welshfella):

y = a(x - 11)^2 - 2 plug in one of the zeroes say y = 0 and x = 10 and solve for a

OpenStudy (anonymous):

0=a(10-11)^2-2 0=-1a a=0

OpenStudy (welshfella):

your secon ine is wrong

OpenStudy (welshfella):

(10-11)^2 = -1^2 = 1

OpenStudy (anonymous):

okay.

OpenStudy (welshfella):

0 = a - 2 a = 2

OpenStudy (welshfella):

so now you need to expand y = 2(x - 11)^2 - 2

OpenStudy (anonymous):

and that is done by..? sorry...i rly dont get this stuff..

OpenStudy (welshfella):

2(x - 11)^2 - 2 = 2[x^2 -22x + 121] - 2 can you continues

OpenStudy (welshfella):

multiply each term in the parentheses by 2

OpenStudy (anonymous):

oh okay i see!

OpenStudy (anonymous):

so now i expand it

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

2(x-11) (x-11)-2

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

no you just multiply 2x^2 - 22x + 121 by 2 then take 2 away

OpenStudy (anonymous):

what? im sorry im confused

OpenStudy (welshfella):

- ive worked out (x - 11)(x - 11) already

OpenStudy (welshfella):

(x - 11)(x - 110 = x^2 - 22x + 121

OpenStudy (welshfella):

* 11) not 110

OpenStudy (anonymous):

okay.. so i multiply the whole thing by 2?

OpenStudy (welshfella):

yes -then take 2 away ( the -2 at the end)

OpenStudy (anonymous):

4x^2-44x+242. now subract 2 from each number?

OpenStudy (welshfella):

no - just from the 242 you can only subtract like terms

OpenStudy (anonymous):

4x2-44+240

OpenStudy (anonymous):

then take away a 2?

OpenStudy (welshfella):

oh its 2*x2 = 2x^2 not 4x^2

OpenStudy (anonymous):

okay i was gonna say. lol everything else matches but the 4x lol

OpenStudy (welshfella):

2x^2 - 44x + 242 - 2 = 2x^2 - 44x + 240

OpenStudy (anonymous):

got it! so i was close, i had a feeling it be that one or the last one.